In a decision tree problem involving capacity planning, there was a random outcome node with three...

  1. In a decision tree problem involving capacity planning, there was a random outcome node with three random outcomes (A,B, and C) which followed a decision node. The probabilities for A and B are P(A)=0.4 and P(B) = 0.5. Payoff following A,B and C respectively are $10000, $5000 and $2000. The expected value of that random outcome node is:
    1. $7400
    2. $6700
    3. $17000
    4. $9800
Answer #1

Expected value of the random outcome mode = P(A)*Payoff of A+P(B)*Payoff of B+P(C)*Payoff of C

P(C) = 1-{P(A)-P(B)

= 10000*0.4+5000*0.5+2000*0.1 = 6700

Option b is right